OpenStudy (anonymous):

line ab contains points A(1,2) and B(-2,6). the slope of line ab is

OpenStudy (anonymous):

try to use : \[m = \frac{ y2-y1 }{ x2-x1 }\]

OpenStudy (jdoe0001):

\(\bf \begin{array}{lllll} &x_1&y_1&x_2&y_2\\ A&({\color{red}{ 1}}\quad ,&{\color{blue}{ 2}})\quad B&({\color{red}{ -2}}\quad ,&{\color{blue}{ 6}}) \end{array} \\\quad \\ slope = {\color{green}{ m}}= \cfrac{rise}{run} \implies \cfrac{{\color{blue}{ y_2}}-{\color{blue}{ y_1}}}{{\color{red}{ x_2}}-{\color{red}{ x_1}}}\)

OpenStudy (anonymous):

so m = -4/3

OpenStudy (anonymous):

(6-2)/(-2-1)

OpenStudy (anonymous):

4/-3